OpenStudy (anonymous):

Question : 4 - Signals: How to check whether the given system is \(\color{green}{\text{Time Invariant}}\) or \(\color{red}{\text{Time Variant}}\)??

OpenStudy (tkhunny):

Perhaps a working example would help?

OpenStudy (tkhunny):

Have you actually SEEN "time-invariance"? In the strictest sense, man made time-invariance doesn't exist. Systems degrade.

OpenStudy (anonymous):

I am just learning the case of LTI Systems, so I want to know the method of checking Time Invariance or Variance of a system..

OpenStudy (anonymous):

Let it be: y(t) = x(-t)...

OpenStudy (tkhunny):

Okay, lag the system and see if you get a different answer. y(t-T) = x(-(t-T)) Is that the same response?

OpenStudy (anonymous):

I did not get what you are trying to say..

OpenStudy (anonymous):

y(t) = x(-t) y(t-T) = x(-(t-T)) = x(-t + T) Now, with what I have to compare it??

OpenStudy (tkhunny):

What does x(-t) mean? Is that 'x' a function or is that 'x' multiplied by "-t"?

OpenStudy (anonymous):

y(t) = output of a system. x(t) = Input given.. x(t) is a signal or x is function on independent variable 't'..

OpenStudy (anonymous):

So, my when I am giving x(t) as an Input, my system is returning x(-t) as Output. :)

OpenStudy (anonymous):

So, my System is designed to do Time-Reversal function..

OpenStudy (tkhunny):

Well, how can Time Reversal be Time-Invariant? The whole question is on whether you get a different response at a different time. If all youg3et is reversal, and that's all it does, I'm not seeing how that is time-variant.

OpenStudy (anonymous):

This you are just saying by manual analyzing.. That's why I said, I know the method to check Time Variance or Invariance of a system.. Any I have never said the system is Time Variant..

OpenStudy (anonymous):

May be you know or not: There is certain procedure to test whether a system is time invariant or not, I want to know that only.. :)

OpenStudy (tkhunny):

We poked at that, earlier. I found your specific problem. http://vimeo.com/6808745

OpenStudy (anonymous):

How to check it for: \[y(t) = \int\limits_{-10}^{10}x(t).dt\]

OpenStudy (anonymous):

Let me try : \[x_1(t) = x_1(t - t_0)\] \[y_1(t) = \int\limits_{-10}^{10}x_1(t) \cdot dt \implies y_1(t) = \int\limits_{-10}^{10}x(t - t_0) \cdot dt\]

OpenStudy (anonymous):

Now, substitute : \(t - t_0 = \lambda \implies dt = d\lambda\) \[y_1(t) = \int\limits_{-10 - t_0}^{10 - t_0} x(\lambda) \cdot d \lambda\]

OpenStudy (anonymous):

While : \[y (t - t_0) = \int\limits_{-10}^{10}x(t - t_0) \cdot dt \color{green}{\implies y_1(t) \ne y(t - t_0)}\] So, given system is \(\color{blue}{\text{Time-Variant..}}\) Am I true in doing that??

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Sorry, spent the day in the hospital... That looks like it! I am a little curious where you got [-10,10]. Is that in the problem statement?

OpenStudy (anonymous):

Sorry, are you alright?? The limits are from -10 to 10 in the question itself..

OpenStudy (tkhunny):

Fair enough. That would have been helpful information up front. :-) Getting better. Thanks.

OpenStudy (anonymous):

I have one doubt left.. In \(y_1(t)\), my integrand is \(x(t)\) And in \(y(t - t_0)\), it is \(x(t - t_0)\), is it right there?